# Towards Fair and Robust Volumetric CT Classification via KL-Regularised Group Distributionally Robust Optimisation

### 16 Mar 2026 v1

@@ Abstract
Automated diagnosis from chest computed tomography (CT) scans faces two persistent challenges in clinical deployment: distribution shift across acquisition sites, and performance disparity across demographic subgroups. We address both simultaneously across two complementary tasks: binary COVID-19 classification from multi-site CT volumes (Task 1) and four-class lung pathology recognition with gender-based fairness constraints (Task 2). Our framework combines a lightweight MobileViT-XXS slice encoder with a two-layer SliceTransformer aggregator for volumetric reasoning, and trains with a KL-regularised Group Distributionally Robust Optimisation (Group DRO) objective that adaptively upweights underperforming acquisition centres and demographic subgroups. Unlike standard Group DRO, the KL penalty prevents group weight collapse, providing a stable balance between worst-case protection and average performance. For Task 2, we define groups at the granularity of genderclass, directly targeting severely underrepresented combinations such as female Squamous cell carcinoma. On Task 1, our best configuration achieves a challenge F1 of 0.835, surpassing the best published challenge entry by . On Task 2, Group DRO with achieves a mean per-gender macro F1 of 0.815, outperforming the best challenge entry by pp and improving Female Squamous F1 by over the Focal Loss baseline.
@@ 1 Introduction
In this work, we address both challenges simultaneously through a group-aware training objective applied to two related tasks: binary COVID-19 classification across four acquisition centres (Task 1), and four-class lung pathology classification with gender-based fairness constraints (Task 2) [12, 11]. We build on Group Distributionally Robust Optimisation (Group DRO) [22], which minimises the worst-case loss over predefined subgroups rather than the average loss, and augment it with a KL regularisation term that prevents the group weights from collapsing onto a single hard group. This KL-regularised Group DRO objective provides a principled mechanism for reweighting underrepresented subgroups such as female Squamous cell carcinoma patients (5 training samples), without discarding information from majority groups.
@@
A volumetric CT classification framework combining a MobileViT-XXS slice encoder with a SliceTransformer aggregator, achieving competitive performance at under 2M parameters on both single-site and multi-centre benchmarks.
@@
A KL-regularised Group DRO training objective that adaptively reweights acquisition centres (Task 1) and fine-grained genderclass subgroups (Task 2), improving worst-case subgroup performance over standard weighted cross-entropy and Focal Loss baselines [20].
@@ 3.1 Overview
Our framework addresses volumetric CT classification under two forms of distribution shift: acquisition-site heterogeneity across four hospitals (Task 1) and demographic imbalance across gender and class combinations (Task 2). The pipeline consists of three components: a slice-level visual encoder, a volumetric aggregator, and a domain-robust training objective. Given a 3D CT volume, the model produces a class prediction of presence or absence of Covid-19 in a single forward pass, without test-time adaptation or site-specific normalisation.
@@ 3.2 Problem Formulation
Let denote a dataset of CT volumes, where , is the class label, and is a group identifier. For Task 1, indexes the four acquisition centres; for Task 2, indexes the subgroups.
@@ 3.3 Model Architecture
Each axial slice is processed independently by a shared MobileViT-XXS backbone [21]. MobileViT-XXS combines lightweight MobileNet convolutions with local Transformer blocks, yielding a strong feature extractor at only 3.42M parameters. The original RGB stem is adapted to single-channel input by averaging the pretrained convolutional weights across the channel dimension, preserving pretrained representations. With the classification head removed, the backbone produces a slice-level embedding , where .
@@ 3.4 Domain-Robust Training
where is the uniform distribution over groups and controls regularisation strength. Setting recovers standard vanilla Group DRO; large forces equal group weighting, recovering empirical risk minimization(ERM).
@@ 4.1 Datasets
This task includes CT volumes collected from four hospitals and medical centres (centres 0-3). The dataset comprises 1,530 volumes (train: 1,222; val: 308; test: held out), with 659 non-COVID and 564 COVID-positive cases in the training split. Each volume is labelled binary (COVID / non-COVID) and annotated with a centre identifier . The challenge metric is the average macro F1 across all four centres:

### 27 Apr 2026 v2

@@ Abstract
Automated diagnosis from chest computed tomography (CT) scans faces two persistent challenges in clinical deployment: distribution shift across acquisition centres, and performance disparity across demographic subgroups. We address both simultaneously across two complementary tasks: binary COVID-19 classification from multi-site CT volumes (Task 1) and four-class lung pathology recognition with gender-based fairness constraints (Task 2). Our framework combines a lightweight MobileViT-XXS slice encoder with a two-layer SliceTransformer aggregator for volumetric reasoning, and trains with a KL-regularised Group Distributionally Robust Optimisation (GDRO) objective that adaptively upweights underperforming acquisition centres and demographic subgroups. Unlike standard GDRO, the KL penalty prevents group weight collapse, providing a stable balance between worst-case protection and average performance. For Task 2, we define groups at the granularity of genderclass, directly targeting severely underrepresented combinations such as female Squamous cell carcinoma. On Task 1, our best configuration achieves a challenge F1 of 0.835, surpassing the best published challenge entry by . On Task 2, GDRO with achieves a mean per-gender macro F1 of 0.815, outperforming the best challenge entry by and improving Female Squamous F1 by over the Focal Loss baseline.
@@ 1 Introduction
In this work, we address both challenges simultaneously through a group-aware training objective applied to two related tasks: binary COVID-19 classification across four acquisition centres (Task 1), and four-class lung pathology classification with gender-based fairness constraints (Task 2) [12, 11]. We build on Group Distributionally Robust Optimisation (Group DRO) [22], which minimises the worst-case loss over predefined subgroups rather than the average loss, and augment it with a KL regularisation term that prevents the group weights from collapsing onto a single hard group. This KL-regularised GDRO objective provides a principled mechanism for reweighting underrepresented subgroups such as female Squamous cell carcinoma patients (5 training samples), without discarding information from majority groups.
@@
A volumetric CT classification framework combining a MobileViT-XXS slice encoder with a SliceTransformer aggregator, achieving competitive performance at under 1/3M parameters on both single-site and multi-centre benchmarks.
@@
A KL-regularised GDRO training objective that adaptively reweights acquisition centres (Task 1) and fine-grained genderclass subgroups (Task 2), improving worst-case subgroup performance over standard weighted cross-entropy and Focal Loss baselines [20].
@@ 3.1 Overview
Our framework addresses volumetric CT classification under two forms of distribution shift: acquisition-site heterogeneity across four hospitals (Task 1) and demographic imbalance across gender and class combinations (Task 2). Figure 1 shows representative CT slices illustrating inter-site variability for Task 1. The pipeline consists of three components: a slice-level visual encoder, a volumetric aggregator, and a domain-robust training objective. Given a 3D CT volume, the model produces a class prediction of presence or absence of Covid-19 in a single forward pass, without test-time adaptation or site-specific normalisation.
@@ 3.2 Problem Formulation
Let denote a dataset of CT volumes, where , is the class label, and is a group identifier, where is the number of slices(64 in our case), H and W are the height and width of each slice, respectively. For Task 1, indexes the four acquisition centres; for Task 2, indexes the subgroups.
@@ 3.3 Model Architecture
Each axial slice is processed independently by a shared MobileViT-XXS backbone [21]. MobileViT-XXS combines lightweight MobileNet convolutions with local Transformer blocks, yielding a strong feature extractor at only 1.3M parameters. The original RGB stem is adapted to single-channel input by averaging the pretrained convolutional weights across the channel dimension, preserving pretrained representations. With the classification head removed, the backbone produces a slice-level embedding , where .
@@ 3.4 Domain-Robust Training
where is the mean cross-entropy loss over all samples in group , is the uniform distribution over groups and controls regularisation strength. Setting recovers standard vanilla Group DRO; large forces equal group weighting, recovering empirical risk minimization(ERM).
@@ 4.1 Datasets
| Centre | Class | Train | Val | Total |
| 0 | COVID-19 | 175 | 43 | 218 |
| Non-COVID-19 | 164 | 45 | 209 | |
| 1 | COVID-19 | 175 | 43 | 218 |
| Non-COVID-19 | 165 | 45 | 210 | |
| 2 | COVID-19 | 39 | 0 | 39 |
| Non-COVID-19 | 165 | 45 | 210 | |
| 3 | COVID-19 | 175 | 42 | 217 |
| Non-COVID-19 | 165 | 45 | 210 | |
| Total | COVID-19 | 564 | 128 | 692 |
| Non-COVID-19 | 659 | 180 | 839 |
@@
This task includes CT volumes collected from four hospitals and medical centres (centres 0-3) as shown in 1. The dataset comprises 1,530 volumes (train: 1,222; val: 308), with 659 non-COVID and 564 COVID-positive cases in the training split. Each volume is labelled binary (COVID / non-COVID) and annotated with a centre identifier . The metric used for this task is the average macro F1 across all four centres: